OpenStudy (anonymous):

Use summation notation to write the sum: 2-[1/2]+[1/8]-...+[1/2048]

OpenStudy (anonymous):

Can you tell me what kind of sum you're facing there? Is it geometric or is it arithmetic?

OpenStudy (anonymous):

It's geometric. I try to use the equation \[a _{n}=a _{1}*r ^{n-1}\] to find the number of terms, but it comes back with an imaginary answer. Not sure what I'm doing wrong there.

OpenStudy (anonymous):

seems alright to me, \(r= -\frac{1}{4}\) right?

OpenStudy (anonymous):

Yeah, but I need to find out the number of terms.

OpenStudy (anonymous):

oh so I think I misunderstood your question, you need to figure out the sum of the expression? Because at the topic it says to write as a sum notation.

OpenStudy (anonymous):

I don't need to find the sum, just write it in sum notation. I need to find the number to put on top of the sigma.

OpenStudy (anonymous):

ok so you need to solve the following equation for the upper bound \[ \large 2 \left( \frac{1}{4} \right)^{n-1}= \frac{1}{2048} \] I would leave the minus away.

OpenStudy (anonymous):

Good idea. I just omitted the negative sign and got the right answer. Thanks!

OpenStudy (anonymous):

you're welcome